You can draw everything you want along this spiral curve. You can draw everything you want along this spiral curve. You can draw everything you want along this spiral curve.

You can draw everything you want along this spiral curve. You can draw everything you want along this spiral curve. You can draw everything you want along this spiral curve.

You can draw everything you want along this spiral curve. You can draw everything you want along this spiral curve. You can draw everything you want along this spiral curve.